# Phase diagram of the one-dimensional three-state Potts model with an additional mean-field interaction

### v1

@@ 3 Results
Our next comment concerns the comparison of the curve for K=0.22K=-0.22, in the right plot of figure 2, and the curve for K=0.19475K=-0.19475, in figure 3. The latter value of $K$ is only slightly larger than that of the point TP1, as can be seen in the phase diagram in figure 1. The comparison shows that at the triple point TP1 there are actually three phases, one with the value of $q$ around $0.2$, one with $q$ around $0.5$, and one with $q$ around $0.8$, and that the phase with $q$ around $0.5$ is not present for $K$ smaller than the value at TP1. The zoom in the right plot of figure 3 has the purpose to show better the jump at the first order transition occurring at the largest temperature. An analogous comparison can be made between the curve for K=0.18K=-0.18 and that for K=0.17K=-0.17, in the two plots of figure 4. These two values of $K$ are somewhat smaller and somewhat larger, respectively, of the value at the triple point TP2. The comparison shows that actually at this point there are again three phases, one with $q$ around $0.3$, one with $q$ around $0.55$, and one with $q=2/3$; the phase with $q$ around $0.55$ is not present for $K$ larger than the value at TP2. As can be guessed from the phase diagram in figure 1, for all values of $K$ larger than that at TP2 the plot of $q$ versus $T$ will be qualitatively similar to that seen at K=0.17K=-0.17, with one phase transition from a value of $q$, smaller than $2/3$, to $q=2/3$. Actually, we find that, for $K$ larger than that at TP2, at the transition the order parameter $q$ always jumps from $1/3$ to $2/3$. In this $K$ range the value of the temperature at the transition increases with $K$ in the way determined in section 3.1.
@@
3.1 The asymptotic behavior of the transition lines for $K\to\pm\infty$
The asymptotic behavior of the transition lines for $K\to\pm\infty$ can be obtained as in the following. We begin by considering the case $K\to+\infty$. In this case we can actually obtain the analytical expression of the whole transition line from the TP2 point to the right, at increasing $K$. For this we note that the numerical computations have shown that the first order transition occuring at this line is always associated to a jump form $q=1/3$ to $q=2/3$. This has suggested to investigate the behavior of the function $\widetilde{\phi}(\beta,K,y)$ at a particular $K$-dependent value of $\beta$, a value such that the function is symmetric with respect the value $y=1/2$. If we find such a value, then, since we have already noticed that the function has a vanishing derivative at $y=2/3$ for any value of $\beta$ and $K$, then for that particular value of $\beta$ the function will have also a vanishing derivative for $y=1/3$. This is a necessary, although not sufficient, condition to have a first-order transition between $y=1/3$ and $y=2/3$. We proceed as follows. Remembering that $a=\exp[\beta(3y-2)]$, after a very simple manipulation we first rewrite the function $\widetilde{\phi}(\beta,K,y)$, defined in Eq. (2), as
@@ 3.1 The asymptotic behavior of the transition lines for $K\to\pm\infty$
giving the analytical expression of the first-order transition line to the right of the point TP2. For values of $K$ smaller than that at TP2, the $K$-dependent value of $\beta$ such that the function $\widetilde{\phi}$ is even with respect to $y=\frac{1}{2}$ is still defined, and the values $y=\frac{1}{3}$ and $y=\frac{2}{3}$ still correspond to points where the value of $\widetilde{\phi}$ is the same and its first derivative with respect to $y$ vanishes; however, they are not global minima in $y$ of $\widetilde{\phi}$. The asymptotic behavior of (25) for large $K$ and $T$ is $K=\frac{T}{2}\ln(2T)$.
@@ 4 Discussion and conclusions
Since the Hamiltonian is not symmetric with respect to $q$, continuous transitions correspond, in the 2D thermodynamic phase diagram, to isolated points; they are associated to codimension-1 singularitites [18]. As a consequence, the phase transition lines in the phase diagram, associated to codimension-0 singularities, correspond to first-order phase transitions. It is interesting also to note that these lines extend to both $K\to+\infty$ and $K\to$-$\infty$. It is especially remarkable that for sufficiently large negative $K$ the equilibrium state does not depend (asymptotically) on $K$, and so also the temperature of the first-order transition does not depend on $K$. This is due to the nature of the interaction of the Potts model, in which, contrary to the usual spin-1 models, the interaction between two spins distinguishes only two situations, i.e., equal or different spin states.
@@
Finally, we mention two interesting avenues for future study. First, it would be highly interesting to investigate the model considered here by varying the spin states beyond the current number of three, to check if our results could be extended. Second, it would be valuable to study the finite-size properties of the model, specifically in connection with determining the Casimir force, thereby generalizing the results presented in Ref. [19].
@@ Appendix A The $m=0$ property of the equilibrium states
Figure 7: Full lines: The quadrupole moment $q$ as a function of the temperature $T$ for fixed values of $K$, i.e. $K=0$ in the left panel and K=0.19475K=-0.19475 in the right panel; the latter is the same plot as in the left panel of figure 3. Dashed lines: the function $q(K,T)$ giving the minimal value of $q$ for which the Ising Hamiltonian (31) has an equilibrium state with $m\neq 0$. Since the function $\widetilde{T}(x)$ is equal to $1$ for $x=0$, for $K=0$ the dashed line in the left panel is simply $q=T$.

### v2

@@ 3 Results
Our next comment concerns the comparison of the curve for K=0.22K=-0.22, in the right plot of figure 2, and the curve for K=0.19475K=-0.19475, in figure 3. The latter value of $K$ is only slightly larger than that of the point TP1, as can be seen in the phase diagram in figure 1. The comparison shows that at the triple point TP1 there are actually three phases, one with the value of $q$ around $0.2$, one with $q$ around $0.5$, and one with $q$ around $0.8$, and that the phase with $q$ around $0.5$ is not present for $K$ smaller than the value at TP1. The zoom in the right plot of figure 3 has the purpose to show better the jump at the first-order transition occurring at the largest temperature. An analogous comparison can be made between the curve for K=0.18K=-0.18 and that for K=0.17K=-0.17, in the two plots of figure 4. These two values of $K$ are somewhat smaller and somewhat larger, respectively, of the value at the triple point TP2. The comparison shows that actually at this point there are again three phases, one with $q$ around $0.3$, one with $q$ around $0.55$, and one with $q=2/3$; the phase with $q$ around $0.55$ is not present for $K$ larger than the value at TP2. As can be guessed from the phase diagram in figure 1, for all values of $K$ larger than that at TP2 the plot of $q$ versus $T$ will be qualitatively similar to that seen at K=0.17K=-0.17, with one phase transition from a value of $q$, smaller than $2/3$, to $q=2/3$. Actually, we find that, for $K$ larger than that at TP2, at the transition the order parameter $q$ always jumps from $1/3$ to $2/3$. In this $K$ range the value of the temperature at the transition increases with $K$ in the way determined in section 3.1.
@@
3.1 The asymptotic behaviour of the transition lines for $K\to\pm\infty$
The asymptotic behaviour of the transition lines for $K\to\pm\infty$ can be obtained as in the following. We begin by considering the case $K\to+\infty$. In this case we can actually obtain the analytical expression of the whole transition line from the TP2 point to the right, at increasing $K$. For this we note that the numerical computations have shown that the first-order transition occurring at this line is always associated to a jump form $q=1/3$ to $q=2/3$. This has suggested to investigate the behaviour of the function $\widetilde{\phi}(\beta,K,y)$ at a particular $K$-dependent value of $\beta$, a value such that the function is symmetric with respect to the value $y=1/2$. If we find such a value, then, since we have already noticed that the function has a vanishing derivative at $y=2/3$ for any value of $\beta$ and $K$, then for that particular value of $\beta$ the function will have also a vanishing derivative for $y=1/3$. This is a necessary, although not sufficient, condition to have a first-order transition between $y=1/3$ and $y=2/3$. We proceed as follows. Remembering that $a=\exp[\beta(3y-2)]$, after a very simple manipulation we first rewrite the function $\widetilde{\phi}(\beta,K,y)$, defined in Eq. (2), as
@@ 3.1 The asymptotic behaviour of the transition lines for $K\to\pm\infty$
giving the analytical expression of the first-order transition line to the right of the point TP2. For values of $K$ smaller than that at TP2, the $K$-dependent value of $\beta$ such that the function $\widetilde{\phi}$ is even with respect to $y=\frac{1}{2}$ is still defined, and the values $y=\frac{1}{3}$ and $y=\frac{2}{3}$ still correspond to points where the value of $\widetilde{\phi}$ is the same and its first derivative with respect to $y$ vanishes; however, they are not global minima in $y$ of $\widetilde{\phi}$. The asymptotic behaviour of (25) for large $K$ and $T$ is $K=\frac{T}{2}\ln(2T)$.
@@ 4 Discussion and conclusions
Since the Hamiltonian is not symmetric with respect to $q$, continuous transitions correspond, in the 2D thermodynamic phase diagram, to isolated points; they are associated to codimension-1 singularities [18]. As a consequence, the phase transition lines in the phase diagram, associated to codimension-0 singularities, correspond to first-order phase transitions. It is interesting also to note that these lines extend to both $K\to+\infty$ and $K\to$-$\infty$. It is especially remarkable that for sufficiently large negative $K$ the equilibrium state does not depend (asymptotically) on $K$, and so also the temperature of the first-order transition does not depend on $K$. This is due to the nature of the interaction of the Potts model, in which, contrary to the usual spin-1 models, the interaction between two spins distinguishes only two situations, i.e., equal or different spin states.
@@
Finally, we mention two interesting avenues for future study. First, it would be highly interesting to investigate the model considered here by varying the spin states beyond the current number of three, to check if our results could be extended. In the Introduction we have already reminded that in the purely mean-field Potts model with $q$ states the transition is first-order for $q\geq 3$ (only in this paragraph, as it is clear from the context, $q$ refers to the number of states of a Potts model, and not to the quadrupole moment). Here we limit to remark that the equilibrium states in these mean-field models are characterized by having q1q-1 spin states equally populated [12]. In this paper we have seen that this particular feature holds, for $q=3$, also for the model with both a mean-field interaction and a nearest-neighbour interaction. This may suggest that this holds also for $q>3$. In other words, the property that the equilibrium states break only partially the symmetry of the Hamiltonian, could be present also for the general $q$-state Potts model with mean-field and nearest-neighbour interactions. The study of this possibility could in principle be performed by extending our analysis that uses both a Hubbard-Stratonovich transformation and a transfer matrix method. A map of the Hamiltonian to a spin model, similarly to what has been done in this work, would be useful to put in evidence the partial symmetry breaking, if present. It can be argued that this map would require a spin model with $q$ states and a considerably more involved representation of the Kronecker delta. As a second possible direction of investigation, it would be valuable to study the finite-size properties of the model, specifically in connection with determining the Casimir force, thereby generalizing the results presented in Ref. [19].
@@ Appendix A The $m=0$ property of the equilibrium states
Figure 7: Full lines: The quadrupole moment $q$ as a function of the temperature $T$ for fixed values of $K$, i.e., $K=0$ in the left panel and K=0.19475K=-0.19475 in the right panel; the latter is the same plot as in the left panel of figure 3. Dashed lines: the function $q(K,T)$ giving the minimal value of $q$ for which the Ising Hamiltonian (31) has an equilibrium state with $m\neq 0$. Since the function $\widetilde{T}(x)$ is equal to $1$ for $x=0$, for $K=0$ the dashed line in the left panel is simply $q=T$.